### v1

# Quantization Effects on Neural Networks Perception: How would quantization change the perceptual field of vision models?

Abstract
Neural network quantization is an essential technique for deploying models on resource-constrained devices. However, its impact on model perceptual fields, particularly regarding class activation maps (CAMs), remains a significant area of investigation. In this study, we explore how quantization alters the spatial recognition ability of the perceptual field of vision models, shedding light on the alignment between CAMs and visual saliency maps across various architectures. Leveraging a dataset of 10,000 images from ImageNet, we rigorously evaluate six diverse foundational CNNs: VGG16, ResNet50, EfficientNet, MobileNet, SqueezeNet, and DenseNet. We uncover nuanced changes in CAMs and their alignment with human visual saliency maps through systematic quantization techniques applied to these models. Our findings reveal the varying sensitivities of different architectures to quantization and underscore its implications for real-world applications in terms of model performance and interpretability. The primary contribution of this work revolves around deepening our understanding of neural network quantization, providing insights crucial for deploying efficient and interpretable models in practical settings.
Index Terms—  Neural Network Quantization, Class Activation Maps (CAM), Model Interpretability, Visual Saliency
1 Introduction
The rapid advancements in deep learning have led to the development of increasingly complex neural network architectures. These models have demonstrated remarkable performance across various tasks, such as image classification, object detection, and segmentation. However, deploying these models on resource-constrained devices remains challenging due to their high computational and memory requirements.
Neural network quantization [1] has emerged as a key technique for addressing the deployment challenges by reducing model size and computational complexity. Quantization involves converting the weights and activations of a deep neural network from floating-point precision to lower bit-width representations. Quantization techniques all evolved from the following equation :
$Q(x)=\text{round}\left(\frac{x}{\Delta}\right)\times\Delta\vspace{-2mm}$ (1)
where $x$ is the continuous value, $Q(x)$ presents the quantized value, and $\Delta$ is the quantization step size, determining the precision of the quantized representation.
By reducing the precision of parameters, quantization enables the delivery of lighter-weight models for efficient execution on devices with limited computational resources, such as smartphones [2], embedded systems [3], and IoT devices [4].
While quantization offers practical benefits, it introduces changes to model behavior that may impact interpretability. Interpretable models are crucial for understanding the decision-making process of neural networks, particularly in safety-critical applications such as medical diagnosis and autonomous driving. Ensuring that quantized models maintain interpretability is essential for building trust and facilitating human-machine collaboration.
Class activation maps (CAMs) [5] are visualization techniques that highlight the regions of an image contributing most to a neural network’s decision. CAMs provide valuable insights into where the model focuses its attention when making predictions, enhancing understanding and interoperability. Some algorithms like Grad-CAM [6] and Grad-CAM++[7] analyze the gradients flowing back into the final convolutional layer of a CNN. These gradients highlight which features in the image contributed most to the model’s decision, and the algorithm uses them to create a CAM pinpointing these important regions. Understanding how quantization affects CAMs can provide insights into the changes in model behavior and decision-making processes.
Visual saliency [8] refers to the perceptual prominence or importance of different regions within an image for the human visual system. Unlike CAMs, which are derived from the internal activations of a neural network, visual saliency is based on human perception and attention. Comparing CAMs with visual saliency maps [8] can help assess the alignment between model attention and human perception and evaluate the interpretability of quantized models.
In this paper, we investigate the impact of quantization on model perception, focusing on CAMs and their alignment with visual saliency maps. We evaluate six diverse neural network architectures on a comprehensive dataset of images from ImageNet [9]. Through rigorous experimentation, we analyze the changes in CAMs induced by quantization and their implications for model interpretability and performance.  [^1] Source Code
Our findings contribute to advancing the understanding of neural network quantization and offer insights crucial for deploying efficient and interpretable models in practical settings.
2 Methodology
Fig. 1: The experimental methodology of experimentation.
In this study, we aim to explore the effects of quantization on neural network model perception. By investigating how quantization affects CAMs across different architectures, we hope to uncover insights that can enhance model efficiency and interpretability in real-world applications.
2.1 Model Selection and Quantization
We selected six widely-used neural network architectures, namely VGG16[10], ResNet50[11], EfficientNet[12], MobileNet[13], SqueezeNet[14], and DenseNet[15], due to their diverse structures and widespread adoption across various computer vision tasks.
Quantization, the focal subject of our study, was performed using the Quantization Aware Training (QAT) module from PyTorch 222 Quantization Pytorch only for inference, in order to simulate adaptive quantization. This technique allows for the quantization of neural network weights and activations while preserving the gradient flow necessary for calculating CAMs. Three Quantization levels were chosen: namely Float 32 (f32), Integer 16 (int16), and Integer 8 (int8). It’s worth noting that other quantization techniques in PyTorch do not support gradient flow during quantization.
2.2 CAM Calculation
CAMs were computed for each image in the subset chosen from the validation set of ImageNet [9] across different quantization levels. CAMs provide insights into the regions of an image that contribute most to a neural network’s decision-making process, facilitating interpretability and understanding of model behavior.
The CamGrad++ [7] method was employed for CAM calculation, leveraging its ability to capture gradient information effectively and produce accurate heatmaps.
2.3 Visual Saliency Maps
Visual Saliency Maps [8] were generated using the SATSal model [16], a state-of-the-art technique for estimating visual saliency that showed promising results for both high and low-level scenes. Visual saliency refers to the perceptual prominence or importance of different regions within an image according to human perception and attention. By comparing CAMs with visual saliency maps, we aim to evaluate the alignment between model attention and human perception across different quantization levels.
2.4 Evaluation Metrics
To assess the similarity or difference between different CAMs at different quantization levels and between CAMs and visual saliency maps, we employed three metrics [17]: Similarity (SIM), Kullback-Leibler Divergence (KLD), and Pearson Correlation (CC).
Similarity (SIM) known as histogram intersection, is used to measure the similarity between two distributions. Introduced in [18] for the task of image matching.
KLD is a general information-theory-based metric. It measures the difference between two probability distributions.
CC, also known as the linear (Pearson) correlation coefficient. It is a metric used for measuring the dependency between two heatmaps. A high correlation between the maps means that the two distributions are closely related and interconnected.
2.5 Experimental Setup
We conducted experiments on a dataset consisting of $10,000$ images from the validation set of ImageNet[9]. Each neural network architecture was quantized at the specified levels using QAT, and CAMs were computed for each image in the dataset. Visual Saliency Maps were generated using the SATSal model for comparison with CAMs.
This comprehensive methodology enabled us to systematically investigate the impact of quantization on model perception and interpretability, providing insights into the behavior of neural networks across different quantization levels and their alignment with human perception.
Model Metric VGG-16 ResNet-50 DenseNet-121 MobileNet-V2 SqueezeNet-1.0 EfficientNet-B0
f32 v. GT SIM $\uparrow$ $0.485\pm 0.093$ $0.455\pm 0.089$ $0.473\pm 0.090$ $\textbf{0.506}\pm\textbf{0.087}$ $0.457\pm 0.090$ $0.487\pm 0.085$
CC $\uparrow$ $0.298\pm 0.227$ $0.260\pm 0.184$ $0.312\pm 0.193$ $\textbf{0.425}\pm\textbf{0.136}$ $0.259\pm 0.249$ $0.335\pm 0.219$
KLD $\downarrow$ $1.038\pm 0.419$ $1.091\pm 0.374$ $0.950\pm 0.363$ $\textbf{0.815}\pm\textbf{0.304}$ $1.057\pm 0.372$ $0.946\pm 0.337$
int16 v. GT SIM $\uparrow$ $0.703\pm 0.271$ $0.717\pm 0.276$ $0.733\pm 0.270$ $\textbf{0.741}\pm\textbf{0.255}$ $0.458\pm 0.091$ $0.717\pm 0.263$
CC $\uparrow$ $0.298\pm 0.229$ $0.255\pm 0.180$ $0.316\pm 0.194$ $\textbf{0.420}\pm\textbf{0.134}$ $0.258\pm 0.248$ $0.322\pm 0.218$
KLD $\downarrow$ $1.038\pm 0.421$ $1.065\pm 0.344$ $0.951\pm 0.373$ $\textbf{0.822}\pm\textbf{0.306}$ $1.054\pm 0.372$ $0.949\pm 0.337$
int8 v. GT SIM $\uparrow$ $0.694\pm 0.269$ $0.658\pm 0.273$ $0.636\pm 0.258$ $\textbf{0.704}\pm\textbf{0.253}$ $0.457\pm 0.091$ $0.679\pm 0.271$
CC $\uparrow$ $0.300\pm 0.231$ $0.247\pm 0.181$ $0.292\pm 0.197$ $\textbf{0.419}\pm\textbf{0.134}$ $0.257\pm 0.249$ $0.270\pm 0.218$
KLD $\downarrow$ $1.031\pm 0.413$ $1.095\pm 0.372$ $0.969\pm 0.371$ $\textbf{0.826}\pm\textbf{0.302}$ $1.054\pm 0.370$ $1.038\pm 0.348$
int16 v. f32 SIM $\uparrow$ $0.926\pm 0.085$ $\textbf{0.964}\pm\textbf{0.068}$ $0.963\pm 0.068$ $0.939\pm 0.062$ $0.926\pm 0.094$ $0.882\pm 0.114$
CC $\uparrow$ $0.878\pm 0.163$ $\textbf{0.939}\pm\textbf{0.168}$ $0.895\pm 0.223$ $0.832\pm 0.099$ $0.915\pm 0.184$ $0.751\pm 0.151$
KLD $\downarrow$ $0.075\pm 0.099$ $0.045\pm 0.125$ $\textbf{0.045}\pm\textbf{0.107}$ $0.050\pm 0.028$ $0.049\pm 0.105$ $0.176\pm 0.097$
int8 v. f32 SIM $\uparrow$ $\textbf{0.921}\pm\textbf{0.087}$ $0.844\pm 0.146$ $0.859\pm 0.111$ $0.916\pm 0.073$ $0.898\pm 0.089$ $0.807\pm 0.167$
CC $\uparrow$ $0.875\pm 0.161$ $0.609\pm 0.285$ $0.560\pm 0.264$ $0.755\pm 0.124$ $\textbf{0.893}\pm\textbf{0.196}$ $0.463\pm 0.235$
KLD $\downarrow$ $0.078\pm 0.098$ $0.289\pm 0.215$ $0.183\pm 0.123$ $0.075\pm 0.036$ $\textbf{0.064}\pm\textbf{0.111}$ $0.396\pm 0.172$
Table 1: Quantitative Results for models: GT represents the human visual saliency map, f32 represents the float 32 precision model, int16 represents the integer 16 precision model, int8 represents the integer 8 precision model
Fig. 2: Qualitative results of CAMs and visual saliency across multiple precisions and models.
3 Results
The obtained results shed light on the alterations in Class Activation Maps (CAMs) across various popular architectures when subjected to quantization. Our analysis revealed nuanced changes in model perception, indicating varying sensitivities of different architectures to quantization. The results of our experimentations are summarized in Table 1.
3.1 Empirical Results
The trend observed across the metrics indicates that int16 precision tends to perform relatively better than both f32 and int8 precision in terms of alignment with ground truth visual saliency maps.
The int16 precision models generally exhibit higher $SIM$ and $CC$ scores with ground truth visual saliency maps compared to both f32 and int8 precision models for most architectures. Additionally, Int 16 precision models often demonstrate lower $KLD$ values, suggesting better preservation of information between CAMs and ground truth visual saliency maps. We hypothesize that this discrepancy may be due to the preservation of noise in the model at higher precision levels, leading to a weaker alignment with human perception. On the other hand, the lower precision of int8 quantization may lead to more information loss or distortion in the CAMs, resulting in reduced alignment with human perception.
We also notice that $MobileNet$ consistently demonstrated robustness to quantization, maintaining high alignment between CAMs and Visual Saliency Maps across different precision levels. $SqueezeNet$ showed a consistent performance across all precision levels in comparison with visual saliency. These consistent results imply that it is less sensitive to quantization and maintains a stable alignment with ground truth visual saliency maps. This robustness could be advantageous in practical applications where maintaining interpretability and model perception is crucial, especially when deploying models on resource-constrained devices.
By establishing a reference point using human visual behavior through comparison with visual saliency maps, we gain valuable insights into how well the model’s attention aligns with human perception. However, it’s equally important to assess the relative changes in network perceptual behavior in comparison to the f32 baseline model.
Models like $ResNet50$ and $DenseNet121$ demonstrate relatively high SIM values, indicating substantial preservation of perceptual information during quantization to int16 precision. This observation suggests that these models maintain a similar visual interpretation capability across different precision levels. In contrast, $EfficientNet$ exhibits the most degradation in all metrics, indicating a comparatively lower degree of preservation of perceptual information under quantization.
The $VGG16$ model displays the highest preservation of similarity between int8 and f32 precision CAMs, indicating its effectiveness in retaining spatial details during quantization. Conversely, $SqueezeNet$ demonstrates the strongest linear correlation and the most aligned probability distributions between int8 and f32 precision CAMs, suggesting its robustness to quantization-induced changes in spatial perception.
These findings suggest that over-parameterized architectures like $VGG16$, $ResNet50$, and $DenseNet121$ exhibit strong robustness to quantization due to their high redundancy in parameters. In contrast, $SqueezeNet$ also demonstrates high robustness, which can be attributed to its efficient architecture and deliberate design choices aimed at maintaining performance under quantization.
3.2 Qualitative Results
Figure 2 depicts an example of an image from the study dataset with the related overlaid saliency map. It also depicts the different CAMs associated with different models and levels of precision.
$VGG-16$, $ResNet50$, and $DenseNet121$ exhibit a low level of impact in terms of changes on CAMs, at the int16 precision. On the other hand, we notice significant shifts when comparing at the int8 quantization level. Which is consistent with the deductions from the empirical results.
Smaller networks like $MobileNet$, and $SqueezeNet$ show fewer changes across lower levels. We can especially notice that CAMs from $SqueezeNet$ remain unchanged.
$EfficientNet$ displays the highest level of changes across the models which means that it is much less robust. This is due to the simplicity of the network architecture and the small number of parameters.
4 Conclusion
This study investigated how quantization affects different neural network architectures using Class Activation Maps (CAMs) to gauge their alignment with human perception. We found that quantization with int16 precision strikes a good balance between efficiency and maintaining human-like perception (measured by SIM, CC, KLD). Notably, MobileNet and SqueezeNet exhibited exceptional robustness due to their efficient designs. Conversely, EfficientNet suffered more under quantization.
These findings offer valuable insights for deploying neural networks on resource-constrained devices. By understanding how different architectures respond to quantization, researchers and practitioners can make informed decisions to achieve a balance between model performance, interpretability, and efficiency in real-world applications.